OpenStudy (anonymous):

Integral sin2x*(1+4cos^2 x)^1/2 dx

ganeshie8 (ganeshie8):

Hint : there is an obvious substitution here

OpenStudy (anonymous):

u = 1+ 4cos^2 x ?

ganeshie8 (ganeshie8):

that should work!

OpenStudy (anonymous):

But when I take the derivitive of that for U substitution I get 8cosxsinx which I am unsure if that helps me...

ganeshie8 (ganeshie8):

recall the trig identity \[ \sin (2x) = 2\sin x \cos x\]

OpenStudy (anonymous):

Thanks!!!!

ganeshie8 (ganeshie8):

\[u = 1+4\cos^2x \\~\\\implies du = 4(2\cos x (-\sin x)) dx = -8 \sin x \cos xdx = -4\sin (2x) dx\\~\\\implies \sin(2x) dx = -\frac{du}{4}\]

ganeshie8 (ganeshie8):

the integral becomes \[\int \sqrt{u} (-\dfrac{du}{4}) = -\frac{1}{4}\int \sqrt{u}~du\]

OpenStudy (anonymous):

-1/6 U^3/2

ganeshie8 (ganeshie8):

looks good! dont forget to substitute back the \(u\)

OpenStudy (anonymous):

May be a little late to ask (; Did I set up the right integral for what this problem was asking?? @ganeshie8

OpenStudy (anonymous):

Also evaluating my integral gives me zero and wolphram has an odd form on the answer...

OpenStudy (alekos):

somethings not right. you obviously shouldn't get 0 as the mass of the wire?

OpenStudy (anonymous):

I am wondering if it is my initial integral set up...

OpenStudy (anonymous):

Or am I just makking a mistake towards the end trying to evaluate it?

ganeshie8 (ganeshie8):

your setup looks good, i see two mistakes in the integrand however..

OpenStudy (alekos):

I got it ds should be \[\sqrt{1+4\cos ^{2}2t}\]

ganeshie8 (ganeshie8):

1) you forgot about absolute valuebars around the density : |sin(2x)| 2) see alekos reply above

OpenStudy (alekos):

so the integral changes somewhat

OpenStudy (anonymous):

would it still be u sub at that point?

OpenStudy (alekos):

don't think that's going to work. let me think

OpenStudy (alekos):

wolfram provides an answer of 5.91577

OpenStudy (alekos):

Is this the answer that you would expect?

OpenStudy (anonymous):

Makes more sense than zero but not sure how to calculate integral without wolphrum.

OpenStudy (alekos):

yes. that is the question. still working on it

OpenStudy (anonymous):

Thanks Alekos. I gotta get some sleep.

OpenStudy (alekos):

what time is it over there?

OpenStudy (alekos):

so according to wolfram we have \[-1/4(\sqrt{2\cos4t+3 })\cos2t - 1/8Ln(2\cos2t + \sqrt{2\cos4t+3})\]

OpenStudy (alekos):

step by step solution attached. in order to get the final result integrate over (0,pi/2) and multiply by 4